# Does Mapo Tofu Contain Coffee? Probing LLMs for Food-related Cultural Knowledge

### v1

@@ F.2 Cultural probing results
Table 8: Probing performance comparison with Russian prompts and FmLAMA-ru sub-dataset. “B/mB” respectively represent abbreviations for BERT and mBERT, “XR” denotes XLM-RoBERTa, “b/l” stands for base/large and “c/u” stands for cased/uncased. Bold and underline represent the best-performing and second-performing cultural group within the same model. The Pearson correlation coefficient between mAP and mWS is 0.70 based on the results above.
Origin Count Bb-ky Bb-kl mB-c mB-u XRb XRl mT5 Avg.
Italy 88 (10.9%) 6.77±2.85 7.01±2.68 1.19±0.15 3.14±2.45 1.90±1.27 2.96±3.44 6.81±1.50 4.25
U.S. 64 (7.9%) 10.72±5.32 10.49±4.97 2.74±1.35 4.84±4.86 3.07±3.62 3.25±3.75 11.65±2.84 6.68
Turkey 19 (2.4%) 3.23±2.85 2.77±0.81 3.58±2.81 4.53±2.97 1.83±0.07 2.37±0.71 1.75±0.39 2.87
Japan 101 (12.5%) 5.56±3.00 5.73±3.43 2.43±1.00 1.94±1.37 3.45±0.68 3.47±0.41 6.77±1.79 4.19
France 67 (8.3%) 2.33±0.74 3.56±1.47 1.40±0.31 2.76±0.17 1.60±0.24 1.62±0.20 2.42±0.12 2.24
U.K. 27 (3.3%) 12.25±6.50 16.39±5.29 4.46±4.18 5.81±5.41 2.11±1.30 3.81±4.73 11.71±4.34 8.08
Mexico 13 (1.6%) 7.40±5.58 4.58±1.71 1.99±0.22 5.58±4.32 1.79±0.17 1.78±0.21 2.28±1.25 3.63
India 32 (4.0%) 4.54±2.64 5.22±1.79 1.20±0.08 2.06±1.10 2.70±3.58 1.68±1.30 5.55±1.09 3.28
Germany 15 (1.9%) 2.26±1.15 3.27±0.71 1.56±0.19 2.74±2.45 2.26±2.37 2.06±1.56 3.51±2.73 2.52
China 54 (6.7%) 6.59±3.41 7.07±3.14 2.10±1.05 2.43±2.01 3.30±3.93 2.80±2.60 6.10±1.59 4.34
Iran 8 (1.0%) 1.74±0.76 8.47±6.64 2.28±0.13 1.50±0.80 1.30±0.48 1.15±0.21 1.56±0.53 2.57
Greece 9 (1.1%) 1.27±0.27 6.00±3.98 2.06±0.09 2.31±0.56 0.97±0.15 1.12±0.11 5.67±4.15 2.77
Spain 21 (2.6%) 2.53±1.12 2.95±1.22 3.84±2.58 2.53±2.12 3.04±2.33 3.87±3.90 4.26±2.11 3.29
Russia 8 (1.0%) 2.04±1.88 3.98±2.35 1.28±0.28 1.89±0.78 3.05±4.75 1.71±1.19 1.82±0.33 2.25
Others 281 (34.8%) 4.46±1.86 5.12±1.67 1.84±0.31 1.93±0.99 1.87±1.11 2.00±1.21 4.24±0.63 3.07
ALL 807 (100.0%) 5.42±2.33 6.09±2.16 2.05±0.71 2.68±1.74 2.31±1.52 2.49±1.70 5.56±0.81 3.80
Table 9: Probing performance comparison on mAP (%) with Korean prompts and FmLAMA-ko sub-dataset. “B/mB” respectively represent abbreviations for BERT and mBERT, “XR” denotes XLM-RoBERTa, “b/l” stands for base/large and “c/u” stands for cased/uncased. Bold and underline represent the best-performing and second-performing cultural groups within the same model.
Origin Count Bb-zh mB-c mB-u XRb XRl mT5 Avg.
Italy 49 (5.8%) 13.30±4.45 10.57±2.58 14.37±2.34 10.52±3.21 11.61±3.41 10.15±1.19 11.75
U.S. 101 (11.9%) 15.92±3.83 14.10±2.67 16.08±1.70 9.00±5.25 9.31±4.08 12.67±0.77 12.85
Turkey 12 (1.4%) 8.26±1.41 4.10±1.61 4.91±1.53 2.43±0.85 2.02±0.54 4.94±0.48 4.44
Japan 114 (13.4%) 11.51±4.77 12.29±2.43 12.89±2.23 7.31±4.70 8.32±4.04 7.85±0.48 10.03
France 70 (8.2%) 15.34±4.47 15.31±3.04 15.29±2.19 10.36±5.22 12.79±3.97 7.90±0.97 12.83
U.K. 38 (4.5%) 23.26±5.66 15.85±4.90 18.49±4.34 14.03±6.94 14.95±4.25 9.00±2.26 15.93
Mexico 19 (2.2%) 17.20±3.14 12.52±1.39 12.63±1.61 10.68±3.44 13.99±1.16 15.19±1.90 13.70
India 24 (2.8%) 5.69±2.14 4.96±2.12 6.79±2.52 3.50±2.36 5.27±1.84 1.01±0.11 4.54
Germany 16 (1.9%) 6.97±2.09 11.28±1.30 13.08±3.16 5.61±4.45 5.74±2.34 7.55±0.64 8.37
China 87 (10.2%) 16.68±7.05 15.47±4.87 17.70±6.07 9.20±5.50 12.71±6.72 8.33±1.34 13.35
Iran 7 (0.8%) 9.25±1.04 5.60±1.05 11.30±6.38 7.38±6.42 8.71±6.74 3.21±0.09 7.58
Greece 7 (0.8%) 4.29±0.79 2.13±0.70 1.93±0.28 2.70±1.17 5.58±2.37 4.35±0.48 3.50
Spain 12 (1.4%) 11.98±1.11 7.52±2.42 10.18±1.07 7.55±2.97 9.46±3.33 11.65±1.92 9.72
Russia 8 (0.9%) 5.48±0.98 2.82±0.84 3.79±0.69 2.56±0.98 3.65±1.62 2.11±0.43 3.40
Others 251 (30.8%) 10.55±3.73 9.90±2.97 10.28±1.60 6.59±3.07 9.38±3.56 5.70±0.41 8.73
ALL 815 (100.0%) 12.99±3.96 11.78±2.65 13.01±2.06 8.05±4.00 9.98±3.69 7.88±0.55 10.62
Table 10: Probing performance comparison with Chinese prompts and FmLAMA-zh sub-dataset. “B/mB” respectively represent abbreviations for BERT and mBERT, “XR” denotes XLM-RoBERTa, “b/l” stands for base/large and “c/u” stands for cased/uncased. Bold and underline represent the best-performing and second-performing cultural groups within the same model.
Origin mB-c mB-u mT5 Average
Italy 9.70±4.24 10.42±6.42 12.61±2.28 10.91
U.S. 14.91±4.71 12.86±5.63 11.45±3.22 13.07
Japan 8.27±3.62 9.04±5.07 6.99±3.13 8.10
France 8.26±3.15 8.13±4.40 5.00±2.18 7.13
U.K. 19.84±7.17 13.97±9.13 13.11±7.97 15.64
India 13.87±4.46 13.94±7.00 8.84±3.44 12.22
China 14.89±3.74 13.23±5.08 9.04±2.73 12.39
Indonesia 8.33±3.68 6.75±2.38 6.79±2.12 7.29
ALL 11.97±3.27 10.98±4.65 9.03±2.77 10.66
(a) EN_en: mAP (%)
Origin mB-c mB-u mT5 Average
Italy 5.77±3.04 6.18±3.57 4.21±1.12 5.39
U.S. 5.11±2.31 5.75±1.41 2.62±0.63 4.49
Japan 10.51±2.87 9.66±4.55 2.75±1.23 7.64
France 5.02±2.41 6.06±1.66 2.76±0.54 4.61
U.K. 3.14±1.06 3.77±0.98 2.93±0.54 3.28
India 14.43±9.56 16.22±7.65 4.05±1.59 11.57
China 3.87±1.39 6.08±3.77 2.85±1.04 4.27
Indonesia 10.60±1.80 8.76±3.04 3.32±0.90 7.56
ALL 6.77±2.26 7.28±1.95 2.99±0.65 5.68
(b) EN_zh: mAP (%)
Origin mB-c mB-u mT5 Average
Italy 7.25±3.72 6.66±5.45 7.15±1.16 7.02
U.S. 13.23±3.64 7.69±5.49 11.42±2.11 10.78
Japan 12.19±3.70 6.45±4.36 7.61±1.81 8.75
France 11.64±4.46 7.25±5.96 7.66±2.84 8.85
U.K. 13.60±6.01 9.17±7.54 11.34±4.36 11.37
India 3.70±1.51 3.03±1.66 1.57±0.18 2.77
China 15.04±3.69 9.09±8.25 8.80±3.81 10.98
Indonesia 9.65±1.93 4.88±2.55 5.41±2.06 6.65
ALL 11.88±3.49 7.22±5.48 8.40±2.22 9.17
(c) ZH_zh: mAP (%)
Origin mB-c mB-u mT5 Average
Italy 1.53±0.20 2.28±0.25 2.52±0.56 2.11
U.S. 2.02±0.21 3.46±0.49 3.45±0.69 2.98
Japan 2.59±0.44 2.74±0.73 2.65±0.15 2.66
France 2.04±0.60 2.41±0.58 2.42±0.44 2.29
U.K. 1.42±0.22 3.40±1.27 2.29±0.35 2.37
India 2.03±0.37 2.53±0.77 2.23±0.91 2.26
China 2.85±0.44 2.99±1.34 2.29±0.56 2.71
Indonesia 2.04±0.27 2.36±0.36 1.25±0.20 1.88
ALL 2.20±0.33 2.85±0.64 2.58±0.34 2.54
(d) ZH_en: mAP (%)
Table 11: Code-switching analysis: Probing results by using prompts without introducing cultural background.
Origin mB-c mB-u mT5 Average
Italy 9.20±4.55 9.97±4.71 11.27±2.42 10.15
U.S. 15.49±4.37 13.05±5.40 11.37±4.10 13.30
Japan 11.81±3.16 11.84±3.99 10.11±3.43 11.25
France 8.29±2.81 8.93±4.29 5.92±1.38 7.71
U.K. 20.02±9.00 15.81±7.65 15.79±6.17 17.21
India 17.95±2.38 21.88±7.42 11.51±4.65 17.11
China 14.77±3.49 13.75±6.35 9.13±3.70 12.55
Indonesia 9.02±1.01 8.75±2.36 8.11±3.19 8.63
ALL 13.06±3.26 12.42±4.46 10.12±3.16 11.87
(a) EN_en: mAP (%)
Origin mB-c mB-u mT5 Average
Italy 7.57±1.67 6.79±1.71 4.18±1.02 6.18
U.S. 6.01±1.71 7.51±1.08 2.86±0.55 5.46
Japan 13.96±2.05 10.45±3.31 4.77±1.61 9.73
France 5.28±1.72 6.63±0.78 2.92±0.70 4.94
U.K. 3.19±0.91 3.60±0.77 4.44±1.47 3.74
India 23.58±2.64 19.51±6.39 6.59±2.59 16.56
China 5.23±0.14 6.39±3.55 3.98±1.08 5.20
Indonesia 10.12±1.01 7.91±2.69 6.35±1.81 8.13
ALL 8.57±1.25 8.09±1.35 4.10±0.98 6.92
(b) EN_zh: mAP (%)
Origin mB-c mB-u mT5 Average
Italy 6.35±3.08 7.67±4.89 6.33±1.80 6.78
U.S. 13.60±3.06 8.50±5.41 10.17±1.91 10.76
Japan 13.02±2.43 7.32±4.47 8.83±1.60 9.72
France 12.12±4.46 9.00±5.67 7.48±2.82 9.53
U.K. 13.97±6.02 8.21±7.81 10.87±4.47 11.02
India 4.60±1.88 3.38±2.06 1.82±0.30 3.27
China 20.49±4.29 10.22±8.87 9.71±4.53 13.47
Indonesia 9.09±3.50 4.12±2.30 5.53±1.74 6.25
ALL 13.07±3.13 7.99±5.34 8.46±2.24 9.84
(c) ZH_zh: mAP (%)
Origin mB-c mB-u mT5 Average
Italy 1.57±0.18 2.73±0.63 2.77±0.61 2.36
U.S. 2.29±0.13 3.61±1.16 4.30±0.73 3.40
Japan 2.79±0.65 2.69±1.35 2.35±0.60 2.61
France 3.02±0.86 2.56±0.70 2.25±0.20 2.61
U.K. 1.77±0.21 2.38±1.43 2.24±0.38 2.13
India 2.30±0.51 2.85±1.29 2.30±0.86 2.48
China 3.12±0.71 2.39±1.70 2.31±0.45 2.61
Indonesia 2.39±0.46 2.28±0.77 1.20±0.12 1.96
ALL 2.54±0.42 2.77±1.07 2.68±0.21 2.66
(d) ZH_en: mAP (%)
Table 12: Code-switching analysis: Probing results by using prompts with introducing cultural background.
Origin Llama2-7b-chat-hf Llama2-13b-chat-hf
EN_en EN_zh ZH_zh ZH_en EN_en EN_zh ZH_zh ZH_en
Italy 42.86 20.41 2.04 2.04 48.98 30.61 6.12 10.20
U.S. 37.62 13.86 4.95 4.95 49.50 27.72 5.94 12.87
Japan 14.91 16.67 3.51 4.39 20.18 15.79 4.39 6.14
France 39.71 22.06 5.88 11.76 48.53 26.47 2.94 10.29
U.K. 36.84 18.42 15.79 5.26 39.47 21.05 15.79 7.89
India 20.83 12.50 4.17 4.17 33.33 12.50 12.50 16.67
China 17.65 23.53 10.59 15.29 21.18 24.71 16.47 15.29
Indonesia 6.90 10.34 0 6.90 24.14 13.79 3.45 3.45
All 27.36 17.91 5.91 7.28 35.04 22.64 7.87 10.43
(a) Performance results of Llama2-chat.
Origin vicuna-7b vicuna-13b
EN_en EN_zh ZH_zh ZH_en EN_en EN_zh ZH_zh ZH_en
Italy 51.02 26.53 8.16 12.24 42.86 30.61 18.37 22.45
U.S. 49.50 27.72 17.82 13.86 37.62 25.74 20.79 22.77
Japan 21.93 14.04 7.02 7.02 16.67 14.04 9.65 7.02
France 44.12 23.53 11.76 13.24 39.71 20.59 14.71 17.65
U.K. 47.37 21.05 15.79 10.53 34.21 31.58 23.68 26.32
India 41.67 4.17 12.50 12.50 16.67 16.67 12.50 12.50
China 23.53 18.82 27.06 15.29 17.65 18.82 20.00 22.35
Indonesia 20.69 6.90 10.34 13.79 20.69 10.34 13.79 13.79
All 36.22 19.69 14.37 12.01 28.15 20.87 16.54 17.72
(b) Performance results of vicuna.
Origin gpt-3.5-turbo
EN_en EN_zh ZH_zh ZH_en
Italy 69.39 48.98 20.41 26.53
U.S. 60.40 39.60 22.77 25.74
Japan 28.95 31.58 10.53 16.67
France 63.24 33.82 17.65 27.94
U.K. 57.89 36.84 26.32 26.32
India 54.17 12.50 12.50 16.67
China 30.59 34.12 31.76 37.65
Indonesia 34.48 13.79 20.69 13.79
All 47.64 34.06 20.28 25.00
(c) Performance results of gpt-3.5-turbo.
Table 13: Accuracy (%) of decoder-only LLMs’ probing results in each cultural group.

### v3

@@ D.1 Probing results in each FmLAMA-la
Table 13: Probing performance comparison with Russian prompts and FmLAMA-ru sub-dataset.
D.2 Language analysis on other multilingual LLMs
Figures 111213, and 14 present a comparison of probing results for the multilingual LLMs—mBERT, mT5, Qwen2, and Llama2—using prompts in various languages on the corresponding filtered sub-datasets. Decoder-only LLMs, such as Qwen2, Llama2, and Llama3, while representing the cutting edge of current LLM technology, exhibit greater performance variation under multilingual prompt settings.
Figure 11: Comparison of probing results on mBERT with prompts in different languages,
Figure 12: Comparison of probing results on mT5 with prompts in different languages,
Figure 13: Comparison of probing results on Qwen2 with prompts in different languages,
Figure 14: Comparison of probing results on Llama2 with prompts in different languages,
D.3 Code-switching probing results in each cultural group
Figure 1528 presents a detailed comparison of the probing results for each cultural group and other multilingual LLMs.
Figure 15: Italy: Probing results of code-switching prompts.
Figure 16: U.S.: Probing results of code-switching prompts.
Figure 17: Turkey: Probing results of code-switching prompts.
Figure 18: Japan: Probing results of code-switching prompts.
Figure 19: France: Probing results of code-switching prompts.
Figure 20: U.K.: Probing results of code-switching prompts.
Figure 21: Mexico: Probing results of code-switching prompts.
Figure 22: India: Probing results of code-switching prompts.
Figure 23: Germany: Probing results of code-switching prompts.
Figure 24: China: Probing results of code-switching prompts.
Figure 25: Greece: Probing results of code-switching prompts.
Figure 26: Spain: Probing results of code-switching prompts.
Figure 27: Russia: Probing results of code-switching prompts.
Figure 28: ALL: Probing results of code-switching prompts.
Appendix E Cultural bias vs Dish popularity
Table 1 shows that some LLMs perform better in probing tasks for low-resource language groups. Specifically, the decoder-only LLM performs well in the Iran group (both mAP and mWS metric), while Greece shows decent results with both T5 and decoder-only LLMs (only mWS metric).
To understand the strong performance of the Iran group across different metrics, we analyzed the results for the 21 dishes. Table 14 shows 17 involved dishes where the predicted ingredients ranked in the top 10 (Llama3). Since the top-ranked predicted ingredients in the table are not all commonly used, we can rule out the possibility that the strong performance of the Iran group is due to the LLM’s common ingredient prediction errors. We hypothesize two possible reasons for this performance: (1) LLM Cultural Ability: Decoder-only LLMs may be more familiar with the Iran cultural group. (2) Data Bias: The dishes in the Iran group are likely well-known and representative of Iranian cuisine, so the LLM might show stronger memory performance for familiar entities Du et al. (2024).
ID Subject (dish) Involved Object (ingredients) Rank
0 pomegranate soup pomegranate 0
1 Albaloo polo rice 1
2 Kofta minced meat 0
3 Sabzi polo herb 5
4 falooda milk, vermicelli 1, 0
5 ghormeh sabzi parsley, herb 0, 4
6 Bastani milk, sugar, egg 8, 7, 6
7 zeytoon parvardeh olive 8
8 rock candy sugar 0
9 Faloodeh vermicelli 0
10 muhallebi milk, rice flour 2, 1
11 rogan josh lamb meat 1
12 ash-e doogh yogurt 0
13 Mahyawa fish 4
14 sholezard rice 1
15 Eggplant Caviar eggplant 0
16 nân-e panjere flour 6
Table 14: Top 17 dishes from the Iran group with ingredients ranked in the top 10 by Llama3, excluding common prediction errors.